OpenStudy (anonymous):

Find the slope of the tangent to the curve y =(1/2x) + 3; at the point where x = -1. Find the angle which this tangent makes with the curve y = 2x^2 + 2

OpenStudy (anonymous):

y = 1/2x + 3 won't have a tangent, since it's linear. Is it supposed to be y = 1/2x^2 + 3?

OpenStudy (anonymous):

I think y=x/2+3 is the tangent line and the curve is y= 2x^2 + 2.

OpenStudy (anonymous):

not satisfied

OpenStudy (anonymous):

Can you clarify the question please?

OpenStudy (anonymous):

@CliffSedge You're right. The question is written quite oddly, so it threw me for what was being asked.